OpenStudy (18jonea):

The range of tides is affected by the relative positions of the sun and the moon. The highest high tides and the lowest low tides occur during the new moon and full moon. During the first and third quarters of the lunar month, the lowest high tides and highest low tides occur. Throughout the month, the tidal range gradually increases and decreases between the minimum and maximum values of the range. 7) Based on the information above, at what time would the difference between the highest y value and the lowest y value of the tide values be the greatest?

OpenStudy (18jonea):

@Michele_Laino

ganeshie8 (ganeshie8):

Hi

OpenStudy (18jonea):

hi @ganeshie8 could you help me?

OpenStudy (18jonea):

@welshfella

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (aveline):

The y-values are the tides, correct? What are the x-values?

OpenStudy (18jonea):

not sure @Aveline

OpenStudy (aveline):

The x-values are the phases of the moon. The tides, or y-values, are dependent on the x-values, or the phases of the moon.

OpenStudy (aveline):

When are the tides highest and lowest?

OpenStudy (18jonea):

so biggest difference is when there is a new moon or full moon?

OpenStudy (18jonea):

@Aveline

OpenStudy (aveline):

You are correct

OpenStudy (18jonea):

Based on the information above, at what time would the difference between the highest y value and the lowest y value of the tide values be the lowest?

OpenStudy (aveline):

"During the first and third quarters of the lunar month, the lowest high tides and highest low tides occur."

OpenStudy (18jonea):

first and third quarters of a lunar month?

OpenStudy (aveline):

Yup :)

OpenStudy (18jonea):

ok i have four more questions that i have all parts figured out but 1 in each of them. could you help with them?

OpenStudy (aveline):

Sure

OpenStudy (aveline):

Sorry I have a test so I'll brb

OpenStudy (18jonea):

give the a) period, b) amplitude, and c) difference between the highest y value and the lowest y value of each of the periodic functions given below. The periods are all either in terms of π or decimal values in these examples, but you may round to the nearest integer if you prefer.

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

i just need this The periods are all either in terms of π or decimal values in these examples, but you may round to the nearest integer if you prefer.

OpenStudy (aveline):

"The period is the distance required for the function to complete one full cycle"

OpenStudy (michele_laino):

for first function, we have to solve this equation, for \(T\) (\(T\) is the unknown period) \[\cos \left( {3\left( {x + T} \right)} \right) = \cos \left( {3x} \right)\]

OpenStudy (18jonea):

i onow but how do i put it in terms of pi

OpenStudy (volleyballlover55):

@Michele_Laino please come back we're not done.. :)

OpenStudy (18jonea):

i thought the period was 2 but that isnt in terms of pi

OpenStudy (aveline):

You don't need to put them in terms of pi I don't think...? "The periods are all either in terms of π or decimal values"?

OpenStudy (michele_laino):

devloping the left side, we get: \[\cos \left( {3x} \right)\cos \left( {3T} \right) - \sin \left( {3x} \right)\sin \left( {3T} \right) = \cos \left( {3x} \right)\]

OpenStudy (michele_laino):

and the solution, is: \[\begin{gathered} \cos \left( {3T} \right) = 1 \hfill \\ \sin \left( {3T} \right) = 0 \hfill \\ \end{gathered} \]

OpenStudy (18jonea):

so would 2 be right as the period for the first one?

OpenStudy (michele_laino):

solving the equation: \[\cos \left( {3T} \right) = 1\] we get, as the littlest solution: \(3T=2 \pi\) please solve for \(T\)

OpenStudy (aveline):

Well, the equation is cos(3x) I would just use 2pi/b or 2pi/3 even though the graph doesn't look the same.

OpenStudy (18jonea):

2.09439510239

OpenStudy (18jonea):

so what would i do to put it in pi terms

OpenStudy (michele_laino):

please in terms of \(\pi\) @18jonea

OpenStudy (aveline):

Don't bother to simplify, just write \[\frac{ 2\pi }{ 3 }\]

OpenStudy (18jonea):

so that would be the period for the first one?

OpenStudy (michele_laino):

if we divide both sides by \(3\), we get: \[\frac{{3T}}{3} = \frac{{2\pi }}{3}\] please simplify

OpenStudy (aveline):

Correct. Can you figure out the second one? Just put 2pi/b

OpenStudy (18jonea):

how do you simplify 2pi/3

OpenStudy (aveline):

You don't. It's already in terms of pi

OpenStudy (18jonea):

@Michele_Laino said to simplify

OpenStudy (michele_laino):

hint: